KyledaGreat:

Ellen has recently inherited $5200, which she wants to deposit into a savings account. She has determined that her two best bets are an account that compounds monthly at an annual rate of 3.3% (Account 1) and an account that compounds daily at an annual rate of 4.8% (Account 2). Step 1 of 2 : Which account would pay Ellen more interest? Account 1 or Account 2

Vocaloid:

technically you really don't need to do any math to solve this (the account that compounds at a higher rate **and** more often gives more total interest) but demonstrating mathematically: A = P(1 + r/n)^(nt) where P is the principal, r is the interest rate as a decimal, n is the # of times compounded per year, t is time in years for account 1: A = P(1 + r/n)^(nt) = 5200(1 + 0.033/12)^(12*t) for account 2: A = 5200(1 + 0.048/365)^(365t) that 365t exponent makes account 2 multiply much faster

KyledaGreat:

ohhh that makes sense

KyledaGreat:

How much would Ellen's balance be from Account 2 over 2.5 years? Round to two decimal places.

Vocaloid:

account 2: A = 5200(1 + 0.048/365)^(365t), where t = 2.5 A = 5200(1 + 0.048/365)^(365*2.5)

KyledaGreat:

5325.94442208 ⋅ 0.5^x^2

Vocaloid:

I would just chuck this whole line into a calculator if you can 5200(1 + 0.048/365)^(365*2.5) = 5862.94

KyledaGreat:

thank you , that's correct

KyledaGreat:

Martha invests $7800 in a new savings account which earns 5.1% annual interest, compounded semi-annually. What will be the value of her investment after 5 years? Round to the nearest cent.

Vocaloid:

A = P(1 + r/n)^(nt) P = 7800 semi-annually --> n = 2 r = 0.051 t = 5 A = 7800(1 + 0.051/2)^(2*5)

KyledaGreat:

41016.3

Vocaloid:

check your input again A = 7800(1 + 0.051/2)^(2*5) A = 7800(1.0255)^(10) = 10033.47

KyledaGreat:

$10033.47 ?

Vocaloid:

yes

KyledaGreat:

thank you